determine the integral from pie/2 to 0 of function sin2x cosx dx


1
Expert's answer
2021-08-15T16:58:50-0400
"\displaystyle\int_{\pi/2}^{0}\sin 2x \cos x dx"

"\int\sin2x \cos x dx=\int2\sin x\cos x \cos xdx"


"u=\cos x, du=-\sin x dx"


"\int2\sin x\cos x \cos xdx=-\int 2u^2du=-\dfrac{2}{3}u^3+C"

"=-\dfrac{2}{3}\cos^3x+C"

"\displaystyle\int_{\pi/2}^{0}\sin 2x \cos x dx=\big[-\dfrac{2}{3}\cos^3x\big]\begin{matrix}
 0 \\
 \pi/2
\end{matrix}=-\dfrac{2}{3}+0"


"\displaystyle\int_{\pi/2}^{0}\sin 2x \cos x dx=-\dfrac{2}{3}"